# From RNNs to Foundation Models: An Empirical Study on Commercial Building Energy Consumption

### v1

@@ Abstract
Accurate short-term energy consumption forecasting for commercial buildings is crucial for smart grid operations. While smart meters and deep learning models enable forecasting using past data from multiple buildings, data heterogeneity from diverse buildings can reduce model performance. The impact of increasing dataset heterogeneity in time series forecasting, while keeping size and model constant, is understudied. We tackle this issue using the ComStock dataset, which provides synthetic energy consumption data for U.S. commercial buildings. Two curated subsets, identical in size and region but differing in building type diversity, are used to assess the performance of various time series forecasting models, including fine-tuned open-source foundation models (FMs). The results show that dataset heterogeneity and model architecture have a greater impact on post-training forecasting performance than the parameter count. Moreover, despite the higher computational cost, fine-tuned FMs demonstrate competitive performance compared to base models trained from scratch.
@@ 1 Introduction
We use the ComStock dataset (Parker et al., 2023) to study the effect of heterogeneity on various STLF models. ComStock is a synthetic dataset containing energy consumption of more than 350,000 U.S. commercial buildings over one year, alongside additional features. These features include building type, allowing us to control heterogeneity in curated ComStock subsets by varying building type diversity. We generate two such datasets, with one exhibiting greater heterogeneity. In this context, we evaluate multiple time series forecasting models from the literature, broadly categorized into two types: base models and foundation models (FMs). Base models are trained from scratch on a given dataset, while FMs are pre-trained on a large corpora, and only need to be fine-tuned for the given downstream application. We sample across a wide cross-section of available base models: recurrent neural networks (RNNs) (LSTM (Hochreiter and Schmidhuber, 1997), LSTNet (Lai et al., 2018)), transformer-based architectures (Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021)), decomposition-based architectures (Autoformer (Wu et al., 2021)), and patch-based architectures (TimesNet (Wu et al., 2023)). For foundation models, we select TimesFM (Das et al., 2024), which handles multivariate inputs and is tested in both zero-shot (ZS) and fine-tuned (FT) modes. This setup enables us to assess the benefits of pre-training alongside the impact of heterogeneity.
This paper makes three key contributions: (i) the introduction of two curated building energy consumption datasets from ComStock, which are of the same size but vary in heterogeneity; (ii) a performance comparison of various base models on these datasets; and (iii) an evaluation of foundation models in both zero-shot and fine-tuned modes, compared to base models.
@@ 2 ComStock Dataset
ComStock, developed by the National Renewable Energy Laboratory and funded by the U.S. Department of Energy, is a detailed model of the U.S. commercial building stock. It uses statistical sampling methods and advanced energy simulations via OpenStudio (Guglielmetti et al., 2011) to estimate energy consumption across more than 350,000 unique building energy models. These models encompases various building characteristics, including building type, construction year, HVAC system, and occupancy schedules, using 15-minute resolution data for one year. The dataset provides detailed outputs like energy consumption across different end uses and critical metrics, resulting in approximately 10 TB of data per release.
For our study, we distill the ComStock data into two curated datasets from buildings in Illinois, termed IL-HET (heterogeneous) and IL-HOM (homogeneous). Each dataset contains 592 buildings for the actual meteorological year 2018. IL-HET includes all 14 available building types in ComStock (see Figure 2(a)), maintaining the overall dataset distribution, thus introducing heterogeneity. In contrast, IL-HOM comprises only buildings of type Warehouse, providing a more homogeneous set due to correlations among similar building loads. The heterogeneity in IL-HET is illustrated through its temporal spread in Figure 2(b), while IL-HOM demonstrates a more uniform pattern (Appendix A).
@@ 3 Experiment Design
where $\mathbf{y}_{t}\in\mathbb{R}^{n}$ is the building energy consumption, $\mathbf{x}\in\mathbb{R}^{m}$ includes the two weather features, $\mathbf{u}_{t}$ denotes two time induces, and $\mathbf{s}$ includes the three static building characteristics. Furthermore, $T$ and $L$ are the ‘lookahead’ and ‘lookback’ durations, respectively. The forecast model $f(\cdot)$ that is parameterized by the vector $\boldsymbol{\theta}\in\mathbb{R}^{h}$ can be trained in a supervised fashion, followed by evaluating its performance on unseen test sets. We use two metrics: mean squared error (MSE) and mean absolute error (MAE), which given time indices of interest $t\in\mathcal{T}$, are defined as $\text{MSE}:=\frac{1}{n|\mathcal{T}|}\sum_{t\in\mathcal{T}}\left\lVert\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\right\rVert_{2}^{2}$ and $\text{MAE}:=\frac{1}{n|\mathcal{T}|}\sum_{t\in\mathcal{T}}\left\lVert\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\right\rVert_{1}$, respectively. While the models are trained using the MSE loss, both MSE and MAE metrics are recorded over the test sets.
The dataset is divided into train, validation, and test sets in a ratio of $0.8:0.1:0.1$ along the time axis for each building. Each set is $z$-normalized using statistics from the train set. Data from all buildings are then concatenated to form unified datasets. Our choice of models (cf. Section 1) aims to provide a comprehensive evaluation of various time series models from the literature. Detailed descriptions of the model categories are provided in Appendix C.
For training the base models and fine-tuning TimesFM, we use the Adam optimizer with default parameters $\beta_{1}=0.9,\beta_{2}=0.999$. Learning rates are tuned per model from the set of $\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5},10^{-5},5\times 10^{-6}\}$, except for Informer and TimesFM. Informer underwent additional tuning on the set $\{10^{-6},5\times 10^{-7},10^{-7}\}$, while TimesFM was also tuned on $10^{-2}$. Base models are trained with large batches of size 1024, and training stops after 20 epochs or if there is no improvement in the MSE metric on the validation set for 5 consecutive epochs.
For fine-tuning TimesFM, we use a smaller batch size of 512 due to its larger memory requirements compared to the base models. The fine-tuning process is performed for 2 epochs, taking a similar amount of time to train each base model for 20 epochs. The zero-shot experiments with TimesFM are performed using the off-the-shelf pre-trained timesfm-1.0-200m (https://huggingface.co/google/timesfm-1.0-200m) without any modification. A detailed description of the model hyperparameters, learning rates, and termination epochs is provided in Appendix D.
@@ 4 Experiment Results
The performance of the trained models on the test set is summarized in Table 1, and predictions on the test set are visualized in Appendix E. Immediately, it is obvious that TimesNet provides the best performance for all lookahead values across both the datasets. This can be attributed to the patch-based architecture of TimesNet, which enables it to utilize the information present in features effectively. The superior performance of LSTM and LSTNet as compared to Transformer-based architectures is also notable, and we hypothesize that recurrent architectures of the former utilize feature information more efficiently than the latter’s tokenization and attention schemes. Furthermore, we clearly see the impact of heterogeneity—all experiments with IL-HOM have lower metrics compared to those with IL-HET except for TimesNet and Autoformer—and even here, the difference is minute. This validates our hypothesis that heterogeneous datasets are more difficult to learn than homogeneous datasets, while demonstrating that architectures such as TimesNet are robust to increasing heterogeneity.
Comparing the base models to FMs, we see that TimesFM when fine-tuned offers performance close to the best base models. This demonstrates the inherent benefit of the vast corpora of information absorbed by the model when it comes to downstream forecasting tasks. On the other hand, the zero-shot performance of TimesFM is inadequate compared to all base models, highlighting the importance of fine-tuning to unlock FMs’ performance. We remark that although we fine-tuned the model for only 2 epochs in this study, further fine-tuning with more epochs could potentially improve accuracy. However, this would require significantly more time and computing resources, as compared to the base models.
We remark that some of the base models terminated much earlier than 20 epochs due to the patience criterion. Specifically, several models terminated training in as few as 1, 2, 5, or 7 epochs; see Appendix D. This early termination highlights the efficiency of these base models in adapting to the datasets, contrasting with the longer fine-tuning process required by TimesFM.
Finally, we analyze the importance of base model size on performance in Figure 2. This not only visually reinforces the impact of increasing heterogeneity on model performance, but also highlights the point that model architecture is more important than brute-forcing model size. Evidently, the trend that relatively small models viz. LSTM, LSTNet, and TimesNet can outperform bigger models highlights the importance of efficiently learning temporal relations and feature-dependence on forecasts, rather than increasing parameter sizes.
Table 1: Experiment results for different models tested on the two datasets.
@@
$(L,T)$ LSTM LSTNet Transformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.1731 0.1517 0.0571 0.1095 0.2725 0.2850 0.1669 0.2858
$(96,48)$ 0.5276 0.3534 0.1463 0.1732 0.4205 0.3443 0.1635 0.2690
$(96,96)$ 0.2154 0.1772 0.0998 0.1420 0.3533 0.3826 0.3065 0.4101
Informer TimesNet TimesFM (ZS) TimesFM (FT)
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.4027 0.4758 0.0340 0.0637 0.5503 0.3093 0.0614 0.0876
$(96,48)$ 0.5053 0.6004 0.0773 0.0938 0.7269 0.3405 0.1021 0.1322
$(96,96)$ 0.5480 0.5861 0.0726 0.0954 0.7583 0.3559 0.1319 0.1462
@@
$(L,T)$ LSTM LSTNet Transformer Autoformer
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.0830 0.1405 0.0542 0.1243 0.1053 0.1765 0.1751 0.2405
$(96,48)$ 0.2619 0.2885 0.1216 0.1918 0.1748 0.2390 0.2123 0.2572
$(96,96)$ 0.1458 0.2008 0.1180 0.1866 0.1601 0.2288 0.2463 0.2802
Informer TimesNet TimesFM (ZS) TimesFM (FT)
MSE MAE MSE MAE MSE MAE MSE MAE
$(96,4)$ 0.0801 0.1598 0.0285 0.0783 0.7463 0.4837 0.0668 0.1047
$(96,48)$ 0.1220 0.1871 0.0960 0.1459 0.7394 0.4794 0.1060 0.1368
$(96,96)$ 0.1387 0.1983 0.1017 0.1534 0.7392 0.4798 0.1286 0.1573
Figure 2: Test set performance of base models for $L=96$ and (a) $T=4$, (b) $T=48$, (c) $T=96$.
@@ 5 Concluding Remarks
In this paper, we tested various base models, as well as a fine-tuned and zero-shot FM on two datasets, one more heterogeneous than the other. The results demonstrate the impact of heterogeneity on model performance, the capability of models like TimesNet to be robust against heterogeneity, the advantage of fine-tuned FMs, and their disadvantage in zero-shot mode. These observations serve towards grounding the problem of heterogeneity in time series models and serve as important stepping stones towards domains such as federated learning, where data heterogeneity is inherent to the domain.
@@ Appendix A Analysis of IL-HET and IL-HOM
Further evidence of the homogeneity of IL-HOM can be inferred from the equivalent figures to Figures 2(a) and 2(b) as presented in Figures 3(a) and 3(b). While both figures contain only one element thanks to IL-HOM only containing buildings of the type Warehouse, the lower spread as well as variance of loads is evident from Figure 3(b) as compared to IL-HET.
@@ Appendix C Model Architectures
Table 6: Model sizes, parameter counts and structural hyperparameters for ($T=96,L=48$).
@@
TimesNet Token size: 128 Encoder layers: 3 $k$ for top-$k$ amplitudes: 5 Num kernels: 3 Inception block intermediate channels: 64 Dropout: 0.1 998,289 3.81 MB
@@
  • Patch-based Architectures: These models split the input sequence into smaller sub-sequences called patches, followed by learning the interactions between different patches to better predict the forecast. In this category, we use TimesNet [Wu et al., 2023], which creates 2D frequency-domain patches from 1D input sequences, followed by passing them through convolution-based Inception blocks [Szegedy et al., 2015] to generate the forecast.
@@
TimesFM [Das et al., 2024] is a decoder-only foundation model built for time series forecasting. The model consists of input layers, stacked transformers, and output layers and considers an MSE training loss function. The pre-trained TimesFM model timesfm-1.0-200m is trained on a vast amount and variety of data with both synthetic data and real-world data. There are two major model hyperparameters, namely, model_dim, which determines the size of processed data, and num_layers, which controls the number of stacked transformers in the model. A summary of the model size, parameter counts, and structural hyperparameters in TimesFM is provided in Table 7.
Table 7: Model sizes, parameter counts and structural hyperparameters.
@@
TimesFM model_dim: $1280$ num_layers: $20$ 200m 854MB

### v2

@@ Abstract
Accurate short-term energy consumption forecasting for commercial buildings is crucial for smart grid operations. While smart meters and deep learning models enable forecasting using past data from multiple buildings, data heterogeneity from diverse buildings can reduce model performance. The impact of increasing dataset heterogeneity in time series forecasting, while keeping size and model constant, is understudied. We tackle this issue using the ComStock dataset, which provides synthetic energy consumption data for U.S. commercial buildings. Two curated subsets, identical in size and region but differing in building type diversity, are used to assess the performance of various time series forecasting models, including finetuned open-source foundation models (FMs). The results show that dataset heterogeneity and model architecture have a greater impact on post-training forecasting performance than the parameter count. Moreover, despite the higher computational cost, finetuned FMs demonstrate competitive performance compared to base models trained from scratch.
@@ 1 Introduction
We use the ComStock dataset (Parker et al., 2023) to study the effect of heterogeneity on various STLF models. It is a synthetic dataset containing energy consumption of the U.S. commercial building stock alongside additional features such as building type. We leverage these features to control heterogeneity in curated ComStock subsets by varying building type diversity. We generate two such datasets, with one exhibiting greater heterogeneity. In this context, we evaluate multiple time series forecasting models from the literature, broadly categorized into two types: base models and foundation models (FMs). Base models are trained from scratch on a given dataset, while FMs are pre-trained on a large corpora, and only need to be finetuned for the given downstream application. We sample across a wide cross-section of available base models: recurrent neural networks (RNNs) (LSTM (Hochreiter and Schmidhuber, 1997), LSTNet (Lai et al., 2018)), transformer-based architectures (Transformer (Vaswani et al., 2017), Informer (Zhou et al., 2021)), decomposition-based architectures (Autoformer (Wu et al., 2021)), patch-based architectures (PatchTST (Nie et al., 2023)), and 2D-backbone architectures (TimesNet (Wu et al., 2023)). For foundation models, we select TimesFM (Das et al., 2024), which handles univariate inputs and is tested in both zero-shot (ZS) and finetuned (FT) modes. This setup enables us to assess the benefits of pre-training alongside the impact of heterogeneity.
This paper makes three key contributions: (i) the introduction of two curated building energy consumption datasets from ComStock, which are of the same size but vary in heterogeneity; (ii) a performance comparison of various base models on these datasets; and (iii) an evaluation of foundation models in both zero-shot and finetuned modes, compared to base models.
@@ 2 ComStock Dataset
ComStock, developed by the National Renewable Energy Laboratory and funded by the U.S. Department of Energy, is a detailed model of the U.S. commercial building stock. It uses statistical sampling methods and advanced energy simulations via OpenStudio (Guglielmetti et al., 2011) to estimate energy consumption across the building stock using 350,000 unique building energy models. These models are created from over 80 input distributions, including building characteristics such as building type, construction year, HVAC system, and occupancy schedules. The dataset provides detailed outputs like energy consumption at 15-minute resolution across different end uses and critical metrics, resulting in approximately 10 TB of data per release.
For our study, we distill the ComStock data into two curated datasets from buildings in Illinois, termed IL-HET (heterogeneous) and IL-HOM (homogeneous). Each dataset contains data from 592 building models for the actual meteorological year 2018. IL-HET includes all 14 available building types in ComStock (see Figure 1(a)), maintaining the overall dataset distribution, thus introducing heterogeneity. In contrast, IL-HOM comprises only buildings of type Warehouse, providing a more homogeneous set due to correlations among similar building loads. The heterogeneity in IL-HET is illustrated through its temporal spread in Figure 1(b), while IL-HOM demonstrates a more uniform pattern (Appendix A). Both datasets are available at at https://huggingface.co/datasets/APPFL/Illinois_load_datasets.
@@ 3 Experiment Design
where, for our experiments, $\mathbf{y}_{t}\in\mathbb{R}^{1}$ is the building energy consumption, $\mathbf{x}\in\mathbb{R}^{2}$ includes the two weather features, $\mathbf{u}_{t}\in\mathbb{R}^{2}$ denotes two time induces, and $\mathbf{s}\in\mathbb{R}^{3}$ includes the three static building characteristics, and $T$ and $L$ are the ‘lookahead’ and ‘lookback’ durations, respectively. The forecast model $f(\cdot)$ that is parameterized by the vector $\boldsymbol{\theta}\in\mathbb{R}^{h}$ can be trained in a supervised fashion, followed by evaluating its performance on unseen test sets. We use two metrics: normalized mean squared error (NMSE) and normalized mean absolute error (NMAE), which given time indices of interest $t\in\mathcal{T}$, are defined as $\text{NMSE}:=\frac{1}{n|\mathcal{T}|}\sum_{t\in\mathcal{T}}\left\lVert\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\right\rVert_{2}^{2}/\sigma_{\mathbf{y}}^{2}$ and $\text{NMAE}:=\frac{1}{n|\mathcal{T}|}\sum_{t\in\mathcal{T}}\left\lVert\hat{\mathbf{y}}_{t}-\mathbf{y}_{t}\right\rVert_{1}/\sigma_{\mathbf{y}}$. Here $\sigma_{\mathbf{y}}$ is the normalizing factor. NMSE is used as the loss function for training.
The dataset is divided into train, validation, and test sets in a ratio of $0.8:0.1:0.1$ along the time axis for each building. Each set is $z$-normalized feature-wise using statistics from the train set, and the standard deviation of energy consumption is used as $\sigma_{\mathbf{y}}$. Data from all buildings are then concatenated to form unified datasets. Our choice of models (cf. Section 1) aims to provide a comprehensive evaluation of various time series models from the literature. Detailed descriptions of the model categories are provided in Appendix C. Since the pretraining of TimesFM was done by Das et al. (2024) for a maximum lookback of 512, this becomes the limiting factor in choosing the lookback of all models. Correspondingly, all models use $L=512$. Furthermore, we only use the energy consumption feature with TimesFM since it is a univariate model.
Owing to the diversity of base models as well as TimesFM, we carry out hyperparameter tuning for inferring the optimal learning rate for each model, heterogeneity level and lookahead separately, the details of which are provided in Appendix D. The base models are trained for 20 epochs, while TimesFM is finetuned using full-parameter fine tuning for 10 epochs. The impact of using parameter-efficient finetuning for TimesFM is in the scope of future work. All models are trained with a batch size of 1024, and the training of each model uses a node with 4 Nvidia A100 GPUs. The final trained model weights are available at https://huggingface.co/APPFL/Building_load_forecasting, while zero-shot experiments with TimesFM are performed using the off-the-shelf pre-trained timesfm-1.0-200m (https://huggingface.co/google/timesfm-1.0-200m) without any modification.
@@ 4 Experiment Results
The performance of the trained models on the test set is summarized in Table 1, and predictions on the test set are visualized in Appendix E. Among the base models, the best performance on all experiments is achieved by TimesNet except one case where PatchTST outperforms it. This implies that the 2D-backbone of TimesNet and the periodicity-based patch generation of PatchTST are both effective in learning the patterns needed to generate high-quality forecasts. On the other hand, the relatively lower performance of Transformer, Autoformer, and Informer can be attributed to loss of temporal information in self-attention mechanisms and embeddings (Zeng et al., 2023). This leads to such models being outperformed by simpler architectures such as LSTM and LSTNet. It is important to note that patch-based architectures such as PatchTST and TimesFM (whose underlying architecture is patch-based) performs better on IL-HET than IL-HOM, which is a phenomenon that could indicate an increase in their performance with higher heterogeneity in the training data. This will be further investigated in future works. We also note that our implementation of early-stopping (i.e. terminating training when continued improvements are no longer observed on the test set) yielded disparate results across IL-HET and IL-HOM on base models, with IL-HET training terminating earlier across most models. This is discussed in Appendix D.
Comparing the base models to TimesFM, we see that the fined-tuned TimesFM is able to outperform the best base models, and by a large margin on the IL-HET dataset. It is also worth pointing out that TimesFM is a univariate model while the best base models, e.g., TimesNet and PatchTST, are multivariate models. This demonstrates the inherent benefit of the vast corpora of information absorbed by the model during pre-training when it comes to downstream forecasting tasks. Nonetheless, we should take the differences in model sizes between these models into account as well. On the other hand, the zero-shot performance of TimesFM is inadequate compared to other base models, highlighting the importance of full-parameter finetuning to unlock FM’s potentials.
Finally, we analyze the importance of base model size on performance in Figure 2. This not only visually reinforces the impact of increasing heterogeneity on model performance, but also highlights the point that model architecture is more important than brute-forcing model size. Evidently, the trend that relatively small models such as LSTM, LSTNet, TimesNet, and PatchTST can outperform bigger models highlights the importance of efficiently learning temporal relations and feature-dependence on forecasts, rather than increasing parameter sizes.
Table 1: Experiment results for different models tested on the two datasets. The best results for each $(L,T)$ pair are in bold, while the second-best is italicized and underlined.
@@
$(L,T)$ LSTM LSTNet Transformer Autoformer Informer
NMSE NMAE NMSE NMAE NMSE NMAE NMSE NMAE NMSE NMAE
$({512},4)$ 0.1059 0.1623 0.0702 0.1280 0.2306 0.2316 0.1040 0.1689 0.5850 0.6808
$({512},48)$ 0.3464 0.2701 0.1512 0.1774 0.2551 0.2447 0.1208 0.1766 0.3248 0.3451
$({512},96)$ 0.1361 0.1568 0.1062 0.1449 0.2384 0.2299 0.1129 0.1727 0.1871 0.1808
TimesNet PatchTST TimesFM (ZS) TimesFM (FT)
NMSE NMAE NMSE NMAE NMSE NMAE NMSE NMAE
$({512},4)$ 0.0289 0.0588 0.0332 0.0598 0.0371 0.0598 0.0078 0.0295
$({512},48)$ 0.0674 0.0924 0.0732 0.0953 0.1072 0.1159 0.0341 0.0687
$({512},96)$ 0.0642 0.0892 0.0641 0.0870 0.1050 0.1125 0.0339 0.0657
@@
$(L,T)$ LSTM LSTNet Transformer Autoformer Informer
NMSE NMAE NMSE NMAE NMSE NMAE NMSE NMAE NMSE NMAE
$({512},4)$ 0.0945 0.1457 0.0557 0.1259 0.1058 0.1572 0.1325 0.1994 0.0920 0.1538
$({512},48)$ 0.3355 0.2858 0.1891 0.2148 0.1707 0.2081 0.2216 0.2762 0.1428 0.1966
$({512},96)$ 0.1326 0.1868 0.1404 0.1997 0.1711 0.2110 0.1831 0.2358 0.1657 0.2264
TimesNet PatchTST TimesFM (ZS) TimesFM (FT)
NMSE NMAE NMSE NMAE NMSE NMAE NMSE NMAE
$({512},4)$ 0.0182 0.0684 0.0432 0.0868 0.0559 0.1025 0.0054 0.0309
$({512},48)$ 0.0572 0.1230 0.1239 0.1864 0.3111 0.2873 0.0417 0.1003
$({512},96)$ 0.0532 0.1163 0.1166 0.1810 0.2911 0.2686 0.0495 0.1065
Figure 2: Test set performance of base models for $L=512$ and (a) $T=4$, (b) $T=48$, (c) $T=96$.
@@ 5 Concluding Remarks
In this paper, we tested various base models, as well as a finetuned and zero-shot FM on two datasets, one more heterogeneous than the other. The results demonstrate the impact of heterogeneity on model performance, the capability of models like TimesNet and PatchTST to be robust against heterogeneity, and the superior capability of finetuned FMs on both datasets. These observations serve towards grounding the problem of heterogeneity in time series models and act as stepping stones in domains such as federated learning, where data heterogeneity is inherent to the domain.
@@ Appendix A Analysis of IL-HET and IL-HOM
Further evidence of the homogeneity of IL-HOM can be inferred from the equivalent figures to Figures 1(a) and 1(b) as presented in Figures 3(a) and 3(b). While both figures contain only one element thanks to IL-HOM only containing buildings of the type Warehouse, the lower spread as well as variance of loads is evident from Figure 3(b) as compared to IL-HET.
@@ Appendix C Model Architectures
Table 6: Model sizes, parameter counts and structural hyperparameters for ($T={512},L=48$).
@@
TimesNet Token size: 128 Encoder layers: 3 $k$ for top-$k$ amplitudes: 5 Num. kernels: 3 Inception block intermediate channels: 64 Dropout: 0.1 998,289 3.81 MB
PatchTST Token size: 128 Num. heads: 3 Encoder layers: 3 Patch length: 16 Stride: 8 FCNN hidden dimension: 512 990,384 3.78MB
@@
  • 2D-backbone architectures: These models convert input 1D temporal sequences into 2D tensors with properties that help the model better identify short- as well as long-term patterns in said sequences. In this category, we use TimesNet [Wu et al., 2023], which uses the Fourier transform to identify periodicity, which is then leveraged for tensor conversion. Following the same, the tensors are passed through 2D convolution-based Inception blocks [Szegedy et al., 2015] to generate the forecast.
  • Patch-based Architectures: These models split the input sequence into smaller sub-sequences called patches, followed by learning the interactions between different patches to better predict the forecast. In this category, we use PatchTST [Nie et al., 2023], which uses patching in addition to channel-independence to generate forecasts.
@@
On the other hand, TimesFM is a decoder-only foundation model built for time series forecasting. The model consists of input layers, stacked transformers, and output layers and considers an NMSE training loss function. The pre-trained TimesFM model timesfm-1.0-200m is trained on a vast amount and variety of data with both synthetic data and real-world data. There are two major model hyperparameters, namely, model_dim, which determines the size of processed data, and num_layers, which controls the number of stacked transformers in the model. A summary of the model size, parameter counts, and structural hyperparameters in TimesFM is provided in Table 7.
Table 7: Model sizes, parameter counts and structural hyperparameters for TimesFM (for a comprehensive list, see [Das et al., 2024])
@@
TimesFM Token size: $1280$ Num. Layers: $20$ 200m 854MB